Date: Fri, 25 Sep 1998 11:36:09 -0700 (PDT)
To: 2mass
Subject: IPAC 2MASS WG Meeting #166 Minutes
Cc: chas, stiening, bgreen
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: MnIS547jGtk0US01L2Z+Cw==

           IPAC 2MASS Working Group Meeting #166 Minutes

Attendees: R.Cutri, G. Kopan, S. Van Dyk, R. Beck, T. Evans, R. Hurt, J. White,
           J. Fowler, H. McCallon, S. Wheelock, R. Tam, D. Kirkpatrick,
           D. Engler, W. Wheaton, J. Mazzarella, T. Jarrett, L. Fullmer,
           B. Nelson


1.) Project Update 
2.) Sampler and Spring Release
3.) Palomar Run
4.) Outreach Update
5.) New Action Items 

     A big turnout, undoubtedly related to the candy brought back from Hawaii.
1.) Project Update 

     Northern OPS are back up again.  The problem of image elongation has been 
greatly improved.  The polar axis alignment has also been improved.  However, 
the pole scans still need to be analyzed.  The backup electronics may be tested

     The K-band jump problem in the south was discussed via a telecon.  R. Hurt
provided a report.  The speculation was that the dome was partially eclipsing 
the aperture.  Every other scan was being affected (the north scans, not the 
south).  What the actual cause may be is stray light spilling over from the 
secondary when the dome moves.  It is hoped that the stray illumination is 
uniform across the array.  Before the dome drive broke, the effect was about 30
DN; after the dome broke, the effect was substantially more, about 600 DN; and,
after the repair, the counts dropped to about 10 DN.  R. Beck and D. Engler will
rerun some individual frame reductions, to see if the effect is uniform across
the frames before coaddition.  In the worst cases seen so far, obvious structure
is apparent in the images, including vignetting; these fields clearly will need
to be rescanned.  Hopefully, the whole problem will just come down to flagging
in QA, since a uniform illumination will not affect source photometry.

     G. Kopan gave an update on the cross-scan photometric bias problem.  A
proposal has been made to the Calibration Working Group for establishment of the
photometric zero point to be done based on the aperture photometry on the cal 
scans.  These are low stellar density fields and therefore difficult to correct
with DMAGCOR.  His tests seem to indicate that the routine corrects well for 
larger-area, higher-density sci scans.  The corrections to cal scans could be 
set as interpolations between adjacent sci scans; this is another possibility.

     On Sept 3 a 15-micron shift was made in the focus set point.  The focus
had been originally set to optimize image quality.  The small shift may actually
help the cross-scan photometric bias problem; the focus curve has a fairly flat
bottom at minimum, so shifting slightly from the point of optimal image quality
does not change the configuration by very much.

     Jupiter was actually scanned recently in the south, but fears were allayed
by a realization by M. Skrutskie that the planet has a m_K ~ -1 (a bright red
star, such as Arcturus, is significantly brighter at K!) and also appears
extended.  So, its effect on adjoining scans should be not as dire as originally

     The QA scan-by-scan webpage has been modified by R. Hurt to work well, and
it is linked to the nightly QA page.  The QA pages will remain on the irsatest 
server, rather than the irsa server.

     One last point discussed by R. Hurt is the "H-band worm" artifact in the
southern data.  This may be either a persistence artifact or due to electronics.
It occurs infrequently, but it will be analyzed further.

2.) Sampler and Spring Release

     The Sampler has been delayed some more.  The plan is know to allow the
External Review Board, Science Team, and selected external scientists to 
"beta-test" the Sampler before the ERB meeting on Oct 30.  The Sampler will then
be released within a few days after the ERB meeting, most likely on or about 
Nov 3.  The catalogs need to be readied for this date.

     The 1999 Spring release was to include all data processed by Oct 1, with
about 3500 tiles, >~38 million point sources, and ~6 million extended sources.
Now, it will more likely consist of ~20 million point sources in the +12 to +18
dec bands with quality scores >~ 7.  This will be a somewhat patchy distribution
of sky within this range.

     The pipeline is currently in good shape, including the southern pipeline
using 2MAPPS 2.1.  (Southern nights that have been processed are to be QA'ed
imminently.)  This means that there will be no more deliveries until v2.2.  The
shift now must be to analysis of the Spring release data.

     To that end, J. Mazzarella provided an IRSA update.  The docs have been
revised, and the catalogs are hidden behind passwords at the moment.  An
artifact overlay is being worked on.  The extended source catalog is ready for
the Sampler, with the point source catalog to be added soon.  R. Hurt had a
comment on the Survey Visualizer, on the difficulty of properly accessing 
regions that are near the edge of the Aitoff projection.  He suggested possibly
being provided with a choice of projections, one of them being a flat 
projection.  J. Mazz would consider this suggestion further.

3.) Palomar Run

     D. Kirkpatrick and B. Nelson recently had 2 nights at the Palomar 200-in.
Only the first night was usable.  D. Kirkpatrick reported that out of 10
candidates, half the stars were late M dwarfs and half were brand-new L dwarfs.
The total count of known L dwarfs is now in the mid-30's.  One of the new
L dwarfs was, in fact, one of B. Nelson's candidates.  Another of the new L
dwarfs (of L3 type) is the nearest case so far, with a K~11 and a proper motion
of 1" per annum.  It appears to have a faint companion, too faint for follow-up
with the Keck LRIS instrument.

     B. Nelson had 16 candidates, 12 of which are new AGNs/QSOs, continuing his
~75% success rate.  Of the other candidates, one is an L dwarf (as pointed out
above); one, a carbon star; one, a galaxy without emission lines; and one, a
M giant.  His count is now 50 red AGNs, found above b=30 degrees.

4.) Outreach Update

     S. Van Dyk is analyzing data taken on 1998 Jul 4 in the south, which 
corresponds to the anti-radiant during the peak of the Leonid meteor shower on 
1998 Nov 17.  The Hubble Space Telescope will be pointed within 1 degree of
this point for ten hours (about 6 orbits).  He will consult with B. Nelson, 
T. Jarrett, D. Kirkpatrick, and R. Hurt to see if there are 2MASS objects of
interest for which we can take advantage of this free time.  Observing plans
are due in the HST Director's office by Sep 30.

     The Austin AAS Meeting in January is looming.  2MASS will once again have
an exhibit there, but this time, with one workstation set-up, in conjunction
with IRSA, which will ideally have a contiguous exhibit.  The object is to have
the workstation to demonstrate the Sampler data interface to meeting 

5.) New Action Items

%                                                      %
&                 ACTION ITEMS                         %
%                                                      %

     No new action items were issued by R. Cutri.